# Energy-Efficient Multi-UAV-Enabled MEC Systems over Space-Air-Ground Integrated Networks

### v2

@@ Abstract
Index Terms:
Space-air-ground integrated network (SAGIN), multiple unmanned aerial vehicles (UAVs), energy-efficient mobile edge computing (MEC), resource allocation, trajectory design.
@@ I Introduction
The advancements in mobile communication networks have led to an exponential increase in the demand for ubiquitous access from mobile users (MUs) [1]. Despite the significant enhancement of terrestrial network convenience afforded by the fifth-generation (5G) technology, connectivity issues, such as weak wireless access capability and unsatisfied computing services, still exist in areas with sparse MUs, due to the limitations in coverage and economic costs [2]. To address these issues and achieve pervasive connectivity in remote regions, the integration of non-terrestrial networks has emerged as a promising strategy in the next-generation communication systems [3] to provide ubiquitous services for the MUs.
@@
Space–air–ground integrated network (SAGIN), as a typical non-terrestrial network, includes three primary components: the space-based network, which includes low Earth orbit (LEO) satellites; the air-based network, encompassing aircraft, unmanned aerial vehicles (UAVs), and high-altitude platforms; and the ground-based network, consisting of terrestrial base stations (BSs), MUs, and Internet of Things (IoT) devices [4]. Satellite-enabled non-terrestrial networks can provide the MUs with extensive connectivity across both territory and time [5]. However, due to the long distance between the satellites and the ground, it faces challenges such as high transmission latency, low transmission rates, and increased energy consumption for long-distance transmissions. In contrast, terrestrial communication performs excellently in urban areas with the high density of MUs, providing high-speed and low-latency services [6]. However, due to its limited coverage, it struggles to meet the communication needs of MUs located in remote regions [7]. To tackle this issue, the UAVs can serve as aerial communication devices to provide communication and computing services. Due to their agile mobility, the UAVs can offer more flexible coverage and more stable communication quality for the MUs [8]. Through this multi-tier connectivity of the SAGIN, the MUs can benefit from the diverse services offered by different service providers including the LEO satellites and the UAVs, thereby the quality of services of all MUs can be improved [9, 10].
@@
Furthermore, the widespread utilization of those famous computation-intensive and latency-sensitive tasks and devices, including the multimedia applications, social networks, and emerging intelligent devices such as wearable devices and smart home controllers, has significantly impacted the MUs. The increase in usage has heightened the demand for efficient communication connectivity and high-speed data transmission. Additionally, it has highlighted the importance of enhancing the processing capabilities of the MUs to effectively manage the growing task loads. As a result, optimizing task processing performance has become a critical issue deserving urgent attention [11, 12]. This requires efficient and reliable communication and computing resources. However, traditional cloud-centric computing models struggle to provide the MUs located at remote areas with high-quality user experiences. To address this issue, mobile edge computing (MEC) has been proposed as a solution that efficiently utilizes computational resources at the network edge [13]. The MEC brings computing and storage resources to the edge of the mobile network, placing them closer to the ground MUs. Through localized data processing and real-time analysis, the energy consumption and latency of task processing can be reduced, and the adaptability of the network is also enhanced [14].
@@
To the best of our knowledge, there is a lack of study on the hierarchical computation offloading in SAGIN that considers the MU-UAV association and resource allocation for multiple UAVs with their controllable flying trajectories, while meeting the latency requirements of the MUs. To fill this research gap, this article investigates an SAGIN-supported MEC system that jointly considers the MU-UAV association, the multiple UAVs trajectory design, transmit power allocation, task offloading decisions, and computing frequency allocation. The main contributions can be summarized as follows:
@@ II Related Works
Currently, some pioneering works began to explore the deployment of SAGIN in next-generation communication networks [15, 16, 17, 18, 19, 20, 21, 22, 23]. Benefited from the global coverage capabilities of the LEO satellites, the SAGIN can provide a new approach for achieving seamless coverage and enhancing the ubiquitous computing services. Specifically, Chen et al. [15] proposed an intelligent user association strategy in a novel SAGIN system, and the capacity and coverage probability are also analysed. Wang et al. [16] exploited the target positioning in a maritime SAGIN system, where the direction-of-arrival of the targets were estimated by the UAVs. Mohamed et al. [17] studied the joint UAV trajectory planning and LEO satellite selection, and the user data rate was maximized. Tang et al. [18] proposed a learning based network traffic control for SAGIN by considering the high mobility of users as well as frequent changing network traffic and link state. Besides, the joint deployment and optimization of the SAGIN and next-generation communication technologies, including the backscatter communications [19], the blockchain-enabled systems [20, 21], the IoT data collections [22], and the reconfigurable intelligent surface-assisted systems [23], have been studied. However, the increasing demand for the computing tasks by the emerging applications from the MUs still cannot be satisfied.
@@
Fortunately, the MEC system, as a new distributed computing paradigm based on wireless communication networks, has attracted widespread attentions. Several pioneer works [24, 25, 26, 27, 28, 29, 30, 31] started to investigate the computing potential of the implementation of the MEC systems towards next-generation wireless communication networks. Ren et al. investigated the resource allocation [24], and the collaboration of cloud computing and edge computing [25] to minimize the latency of all mobile devices. Cao et al. [26] proposed a new cooperation method in both computation and communication resources for MEC systems to improve the energy efficiency. Wang et al. [27] studied a wireless powered MEC system, where the total transmission energy consumption was minimized considering both energy and task causality constraints. Zhang et al. [28] optimized the sum of utilities, which jointly considered the energy efficiency, time latency, and price of offloading computations in an intelligent reflecting surface aided MEC system. Liang et al. [29] studied the joint task offloading, communication, and computation resource allocation in a multi-user MEC system. Meanwhile, the UAV-enabled MEC systems were studied in [30, 31], where the systems are benefited by the mobility of the UAV through optimizing the UAV trajectories and resource allocations. Nevertheless, these works mainly focused on the terrestrial networks. The MUs in remote areas are still unable to utilize cloud and edge computing resources to handle computational tasks due to the lack of effective network access.
@@
Introducing the SAGIN into the MEC systems has become a key emerging approach to address the aforementioned issues. Notably, the LEO satellites provide extensive global coverage, while the UAVs offer flexible deployment and dynamic response capabilities. This combination not only compensates for the limitations of single networks but also delivers more stable and efficient computing and communication services across various application scenarios, significantly enhancing task processing efficiency. There has been several works dedicated to developing strategies for efficient task processing and computing resource management in SAGIN-supported MEC systems [32, 33, 34, 35, 36, 37, 38]. Specifically, Huang et al. [32] studied the minimization of total energy consumption for task processing in a SAGIN-supported MEC system. Hu et al. [33] focused on the maximization of total energy efficiency by jointly optimizing the trajectories of two UAVs with different operations, computing resource allocation, and bandwidth allocation. Nguyen et al. [34] studied the computation offloading problem in hybrid edge-cloud based SAGIN, where the total system energy consumption was minimized. Du et al. [35] presented an MEC and blockchain enabled SAGIN architecture, while minimized the energy consumption through the task segmentation, the UAVs and satellite’s bandwidth allocation among their served IoT devices. Zhu et al. [36] jointly optimized system energy consumption and delay costs in an SAGIN framework by designing the resource allocation, the task offloading, and the channel allocation. Li et al. [37] studied the offloading optimization and bandwidth allocation in a joint UAV-LEO satellite offloading SAGIN. Xie et al. [38] aimed to optimize the delay and fair utility of the SAGIN and improve the utilization of the UAVs and the satellites.
@@
However, existing works critically neglect four fundamental challenges in multi-UAV-enabled MEC systems over SAGIN: (1) the absence of hybrid access mechanism optimization for enhanced spectral efficiency; (2) insufficient joint coordination in hierarchical computation offloading and scheduling (e.g., LEO satellite and UAVs), and the MUs’ device-side processing, to fully utilize the computation capability; (3) the absence of energy-efficiency-aware trajectory design across multiple UAVs for flight energy conservation; (4) unexplored flexible channel-aware dynamic association strategies between UAVs and MUs to enable precision offloading decisions. These unresolved limitations severely constrain energy efficiency improvement for the whole SAGIN-supported MEC system.
@@ III System model and Problem Formulation
As depicted in Fig. 1, we consider an SAGIN-supported MEC system, which consists of $M$ MUs, denoted as =$\mathcal{M}=\{1,2,\cdots,M\}$, $K$ UAVs, denoted as $\mathcal{K}=\{1,2,\cdots,K\}$, and an LEO satellite. The LEO satellite and each UAV are equipped with an onboard computing server, which can provide computing services to the MUs. Meanwhile, each MU has a computation task, which can be separated into several independent and fine-grained subtasks [39]. Due to the computing latency requirements of the tasks and the limited computational capabilities of the MUs, a portion of these tasks need to be offloaded to the UAVs and the LEO satellite for remote computing.
We consider a time period $\mathcal{T}$ with $N$ time slots, denoted as $\mathcal{N}=\{1,2,\cdots,N\}$, with each time slot having equal duration $\tau$. At the beginning of each time slot, a task will be generated at each MU. The hybrid time division and frequency division multiple access is considered [40]. Since the duration $\tau$ is very small, the position of the UAV during each time slot is fixed, while varies across different time slots. It is assumed that the UAVs fly at a fixed altitude $H$, which enables them to fly smoothly and avoid obstacles and buildings. Hence, the position of the UAV $k$ can be denoted as ${\mathcal{Q}}_{k}[n]=(\boldsymbol{q}_{k}[n],H)^{\mathsf{}}$, where $\boldsymbol{q}_{k}[n]=(x_{k}[n],y_{k}[n])^{\mathsf{}}$ is the horizontal coordinates. Moreover, the location of the MU $m$ is denoted as ${\mathcal{S}}_{m}[n]=(\boldsymbol{s}_{m}[n],0)^{\mathsf{}}$, where $\boldsymbol{s}_{m}[n]=(x_{m}[n],y_{m}[n])^{\mathsf{}}$.
@@ III-A UAV Kinetic Model
$\|\boldsymbol{v}_{k}[n]\|=\frac{\|\boldsymbol{q}_{k}[n]-\boldsymbol{q}_{k}[n-1]\|}{\tau}\leq V_{\max},\ \forall k,\forall n\in\mathcal{N}\backslash\{1\},$ (2)
where $d_{\min}$ denotes the minimum safe distance between any two UAVs, and $V_{\max}$ denotes the maximum speed of the UAVs. Furthermore, the energy consumption for the flight of the UAV $k$ can be expressed as [41, 42, 43]
@@
where $\zeta_{1}$ and $\zeta_{2}$ are fixed parameters related to the UAV’s weight, wing area, air density, etc [44].
@@ III-B Communication Model
We assume that each MU is equipped with two communication interfaces: one for communicating with the LEO satellite and another for communicating with the associated UAV. The two interfaces operate on distinct spectral bands, ensuring that there is no interference between the communications with the UAVs and the communications with the LEO satellite [45, 46].
Due to the flexibility of the UAVs, they can fly close to MUs to establish line-of-sight (LoS) links. The channel gain between the MU $m$ and UAV $k$ can be expressed as
@@
Besides, a portion of the task of MUs can be offloaded to the LEO satellite for computing. Accordingly, the channel conditions between the MUs and the LEO satellite are primarily influenced by the distance and meteorological environment. We assume that the meteorological environment remains constant during the task processing period, hence the channel gain between each MU and the LEO satellite is predominantly determined by the satellite’s altitude[11, 47, 46]. Thus, the transmission rate between the MU $m$ and the LEO satellite can be expressed as

### v3

@@ I INTRODUCTION
The advancements in mobile communication networks have led to an exponential increase in the demand for ubiquitous access from mobile users (MUs) [Liu2018Space-Air-Ground]. Despite the significant enhancement of terrestrial network convenience afforded by the fifth-generation (5G) technology, connectivity issues, such as weak wireless access capability and unsatisfied computing services, still exist in areas with sparse MUs, due to the limitations in coverage and economic costs [Zhang2023Multiagent]. To address these issues and achieve pervasive connectivity in remote regions, the integration of non-terrestrial networks has emerged as a promising strategy in the next-generation communication systems [Wang2021Survey, 10980172] to provide ubiquitous services for the MUs.
@@
Space–air–ground integrated network (SAGIN), as a typical non-terrestrial network, includes three primary components: the space-based network, which includes low Earth orbit (LEO) satellites; the air-based network, encompassing aircraft, unmanned aerial vehicles (UAVs), and high-altitude platforms; and the ground-based network, consisting of terrestrial base stations (BSs), MUs, and Internet of Things (IoT) devices [Liu2020Task-Oriented]. Satellite-enabled non-terrestrial networks can provide the MUs with extensive connectivity across both territory and time [9508471]. However, due to the long distance between the satellites and the ground, it faces challenges such as high transmission latency, low transmission rates, and increased energy consumption for long-distance transmissions. In contrast, terrestrial communication performs excellently in urban areas with the high density of MUs, providing high-speed and low-latency services [7901477]. However, due to its limited coverage, it struggles to meet the communication needs of MUs located in remote regions [Wang2024Hybrid]. To tackle this issue, the UAVs can serve as aerial communication devices to provide communication and computing services. Due to their agile mobility, the UAVs can offer more flexible coverage and more stable communication quality for the MUs [Wu2018Joint, 10972043]. Through this multi-tier connectivity of the SAGIN, the MUs can benefit from the diverse services offered by different service providers including the LEO satellites and the UAVs, thereby the quality of services of all MUs can be improved [Dazhi2022Terminal-Aware, Bakambekova2024On].
@@
Furthermore, the widespread utilization of those famous computation-intensive and latency-sensitive tasks and devices, including the multimedia applications, social networks, and emerging intelligent devices such as wearable devices and smart home controllers, has significantly impacted the MUs. The increase in usage has heightened the demand for efficient communication connectivity and high-speed data transmission. Additionally, it has highlighted the importance of enhancing the processing capabilities of the MUs to effectively manage the growing task loads. As a result, optimizing task processing performance has become a critical issue deserving urgent attention [Liu2023Energy-Efficient, Zhang2024A]. This requires efficient and reliable communication and computing resources. However, traditional cloud-centric computing models struggle to provide the MUs located at remote areas with high-quality user experiences. To address this issue, mobile edge computing (MEC) has been proposed as a solution that efficiently utilizes computational resources at the network edge [Mao2017ASurvey]. The MEC brings computing and storage resources to the edge of the mobile network, placing them closer to the ground MUs. Through localized data processing and real-time analysis, the energy consumption and latency of task processing can be reduced, and the adaptability of the network is also enhanced [Mach2017Mobile].
@@
To the best of our knowledge, there is a lack of study on the hierarchical computation offloading in SAGIN that considers the MU-UAV association and resource allocation for multiple UAVs with their controllable flying trajectories, while meeting the latency requirements of the MUs. To fill this research gap, this article investigates an SAGIN-supported MEC system that jointly considers the MU-UAV association, the multiple UAVs trajectory design (TD), transmit power allocation, task offloading decisions, and computing frequency allocation. The main contributions can be summarized as follows:
@@ II RELATED WORKS
Currently, some pioneering works began to explore the deployment of SAGIN in next-generation communication networks [9953964, 9380358, 10458883, 9606690, 9963692, 9918062, 10233456, 10454605, 9964037, 11153428, 11153721]. Benefited from the global coverage capabilities of the LEO satellites, the SAGIN can provide a new approach for achieving seamless coverage and enhancing the ubiquitous computing services. Specifically, Chen et al. [9953964] proposed an intelligent user association strategy in a novel SAGIN system, and the capacity and coverage probability are also analysed. Wang et al. [9380358] exploited the target positioning in a maritime SAGIN system, where the direction-of-arrival of the targets were estimated by the UAVs. Mohamed et al. [10458883] studied the joint UAV trajectory planning and LEO satellite selection, and the user data rate was maximized. Tang et al. [9606690] proposed a learning based network traffic control for SAGIN by considering the high mobility of users as well as frequent changing network traffic and link state. Besides, the joint deployment and optimization of the SAGIN and next-generation communication technologies, including the backscatter communications [9963692], the blockchain-enabled systems [9918062, 10233456], the IoT data collections [10454605], and the reconfigurable intelligent surface-assisted systems [9964037], have been studied. Xu et al. [11153428] investigated the energy-efficient optimization problem in an SAGIN system for vehicular communications. Han et al. [11153721] studied the data freshness optimization problem for the wireless power tranfer-assisted data collection system over the SAGIN. However, the increasing demand for the computing tasks by the emerging applications from the MUs still cannot be satisfied.
@@
Fortunately, the MEC system, as a new distributed computing paradigm based on wireless communication networks, has attracted widespread attentions. Several pioneer works [8387798, 8664595, 8488502, 8960510, 9417469, 10373153, 10606316, 10417719] started to investigate the computing potential of the implementation of the MEC systems towards next-generation wireless communication networks. Ren et al. investigated the resource allocation [8387798], and the collaboration of cloud computing and edge computing [8664595] to minimize the latency of all mobile devices. Cao et al. [8488502] proposed a new cooperation method in both computation and communication resources for MEC systems to improve the energy efficiency. Wang et al. [8960510] studied a wireless powered MEC system, where the total transmission energy consumption was minimized considering both energy and task causality constraints. Li et al. [8876867] discussed the online incentive mechanism design integrating computation and communication resource allocation for collaborative task offloading in an MEC system. Yi et al. [8606230] investigated the multi-user task offloading and scheduling considering tradeoffs between the local computing and the edge computing for a multi-access MEC system. Zhang et al. [9417469] optimized the sum of utilities, which jointly considered the energy efficiency, time latency, and price of offloading computations in an intelligent reflecting surface aided MEC system. Liang et al. [10373153] studied the joint task offloading, communication, and computation resource allocation in a multi-user MEC system. Meanwhile, the UAV-enabled MEC systems were studied in [10606316, 10417719], where the systems are benefited by the mobility of the UAV through optimizing the UAV trajectories and resource allocations. Nevertheless, these works mainly focused on the terrestrial networks. The MUs in remote areas are still unable to utilize cloud and edge computing resources to handle computational tasks due to the lack of effective network access.
@@
Introducing the SAGIN into the MEC systems has become a key emerging approach to address the aforementioned issues. Notably, the LEO satellites provide extensive global coverage, while the UAVs offer flexible deployment and dynamic response capabilities. This combination not only compensates for the limitations of single networks but also delivers more stable and efficient computing and communication services across various application scenarios, significantly enhancing task processing efficiency. There has been several works dedicated to developing strategies for efficient task processing and computing resource management in SAGIN-supported MEC systems [10440193, Hu2023Joint, 10342725, 10579794, 10879508, 10891825, 10947633, 11134095, 11145097, 11192086, 11124243]. Specifically, Huang et al. [10440193] studied the minimization of total energy consumption for task processing in a SAGIN-supported MEC system. Hu et al. [Hu2023Joint] focused on the maximization of total energy efficiency by jointly optimizing the trajectories of two UAVs with different operations, computing resource allocation, and bandwidth allocation. Nguyen et al. [10342725] studied the computation offloading problem in hybrid edge-cloud based SAGIN, where the total system energy consumption was minimized. Du et al. [10579794] presented an MEC and blockchain enabled SAGIN architecture, while minimized the energy consumption through the task segmentation, the UAVs and satellite’s bandwidth allocation among their served IoT devices. Zhu et al. [10879508] jointly optimized system energy consumption and delay costs in an SAGIN framework by designing the resource allocation, the task offloading, and the channel allocation. Li et al. [10891825] studied the offloading optimization and bandwidth allocation in a joint UAV-LEO satellite offloading SAGIN. Xie et al. [10947633] aimed to optimize the delay and fair utility of the SAGIN and improve the utilization of the UAVs and the satellites. Zhang et al. [11134095] studied the UAV deployment and the computation offloading optimization for the energy-aware SAGIN. Hsu et al. [11145097] proposed a service caching and task offloading scheme for the MEC service optimization with SAGIN. Liu et al. [11192086] investigated the latency minimization problem for the IoT system over SAGINs. From a task offloading ratio perspective, Huynh et al. [11124243] optimized the bandwidth allocations and the task offloading portions to improve the quality of the MEC services.
@@
However, existing works critically neglect four fundamental challenges in multi-UAV-enabled MEC systems over SAGIN: (1) the absence of hybrid access mechanism optimization for enhanced spectral efficiency; (2) insufficient joint resource coordination in hierarchical communication and computation optimization (HCCO), scheduling (e.g., LEO satellite and UAVs), and the device-side local processing of MUs, to fully utilize the computation capability; (3) the absence of energy-efficiency optimization (EEO) with TD across multiple UAVs for flight energy conservation; (4) unexplored flexible channel-aware dynamic association strategies (DASs) between UAVs and MUs to enable precision offloading decisions. These unresolved challenges constrain energy efficiency improvement for the whole SAGIN-supported MEC system. To clearly illustrate our contributions, Table I summarizes the key differences between our work and prior works related to the SAGIN-supported MEC systems, particularly in terms of the system design and the optimization objective.
TABLE I: Comparison of Existing Works
Works Hybrid Access Multiple UAVs TD DAS EEO HCCO
[10440193, 10342725] $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\times$
[Hu2023Joint] $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\times$
[10579794, 10879508, 10947633] $\times$ $\surd$ $\times$ $\times$ $\times$ $\times$
[10891825] $\times$ $\surd$ $\surd$ $\times$ $\surd$ $\times$
[11134095, 11145097, 11124243] $\times$ $\surd$ $\times$ $\surd$ $\times$ $\times$
[11192086] $\times$ $\surd$ $\surd$ $\times$ $\times$ $\times$
[…]
@@ III SYSTEM MODEL AND PROBLEM FORMULATION
As depicted in Fig. 1, we consider an SAGIN-supported MEC system, which consists of $M$ MUs, denoted as $\mathcal{M}=\{1,2,\cdots,M\}$, $K$ UAVs, denoted as $\mathcal{K}=\{1,2,\cdots,K\}$, and an LEO satellite. The LEO satellite and each UAV are equipped with an onboard computing server, which can provide computing services to the MUs. Meanwhile, each MU has a computation task, which can be separated into several independent and fine-grained subtasks [Zeng2023Joint]. Due to the computing latency requirements of the tasks and the limited computational capabilities of the MUs, a portion of these tasks need to be offloaded to the UAVs and the LEO satellite for remote computing. [^1] A promising future direction lies in integrating digital twin (DT) and AI technologies with the MEC systems over the SAGINs [10238695]. DTs can create dynamic virtual replicas of the SAGINs, enabling predictive resource management and intelligent orchestration for heterogeneous tasks [10234396, 10540318], particularly in emerging applications such as personalized healthcare and human-virtual interaction systems.
We consider a time period $\mathcal{T}$ with $N$ time slots, denoted as $\mathcal{N}=\{1,2,\cdots,N\}$, with each time slot having equal duration $\tau$. At the beginning of each time slot, a task will be generated at each MU. The hybrid time division and frequency division multiple access is considered [Nguyen2022Joint]. Since the duration $\tau$ is very small, the position of the UAV during each time slot is fixed, while varies across different time slots. It is assumed that the UAVs fly at a fixed altitude $H$, which enables them to fly smoothly and avoid obstacles and buildings. Hence, the position of the UAV $k$ can be denoted as ${\mathcal{Q}}_{k}[n]=(\boldsymbol{q}_{k}[n],H)^{\mathsf{}}$, where $\boldsymbol{q}_{k}[n]=(x_{k}[n],y_{k}[n])^{\mathsf{}}$ is the horizontal coordinates. Moreover, the location of the MU $m$ is denoted as ${\mathcal{S}}_{m}[n]=(\boldsymbol{s}_{m}[n],0)^{\mathsf{}}$, where $\boldsymbol{s}_{m}[n]=(x_{m}[n],y_{m}[n])^{\mathsf{}}$.
@@ III-A UAV KINETIC MODEL
[…]
where $d_{\min}$ denotes the minimum safe distance between any two UAVs, and $V_{\max}$ denotes the maximum speed of the UAVs. Furthermore, the energy consumption for the flight of the UAV $k$ can be expressed as
@@
where $\zeta_{1}$ and $\zeta_{2}$ are fixed parameters related to the UAV’s weight, wing area, air density, etc [Zeng2017Energy-Efficient].
@@ III-B COMMUNICATION MODEL
We assume that each MU is equipped with two communication interfaces: one for communicating with the LEO satellite and another for communicating with the associated UAV. The two interfaces operate on distinct spectral bands, ensuring that there is no interference between the communications with the UAVs and the communications with the LEO satellite [Zhang2014Dynamic, Zhou2021Deep].
Due to the flexibility of the UAVs, they can fly close to MUs to establish line-of-sight (LoS) links [Diao2020Fairness-Aware, Zeng2017Energy-Efficient, Nguyen2022Joint, Wu2018Joint, 10606316]. The channel gain between the MU $m$ and UAV $k$ can be expressed as
@@
Besides, a portion of the task of MUs can be offloaded to the LEO satellite for computing. Accordingly, the channel conditions between the MUs and the LEO satellite are primarily influenced by the distance and meteorological environment. We assume that the meteorological environment remains constant during the task processing period, hence the channel gain between each MU and the LEO satellite is predominantly determined by the satellite’s altitude[Liu2023Energy-Efficient, Zhang2023Learning, Zhou2021Deep]. Thus, the transmission rate between the MU $m$ and the LEO satellite can be expressed as